An object starts from rest with constant accelration 26m/s^2 a long a linear line. Find the distance travelled in 10th second?

1
Expert's answer
2022-05-15T17:03:20-0400
"d=v_0t+\frac{at^2}{2}"

"d=\rm 0\: m/s\times 10\: s+\frac{26\: m/s^2\times (10\: s)^2}{2}
=1300\: m"